OpenStudy (anonymous):

Please help! A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will the height of the ball be after 12 bounces? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

You dont have explain in sentences. I just want the height please!

OpenStudy (anonymous):

This is geometric sequence

OpenStudy (anonymous):

Yes it is, I get the answer 0.48828125, But i dont think it is right.

OpenStudy (anonymous):

a = 1000 r = 1/2 n = 12 \[1000 (\frac{1}{2})^{n-1}\]

OpenStudy (anonymous):

It should be n not n-1

OpenStudy (anonymous):

Ok I did that and get 0.488281125

OpenStudy (anonymous):

\[1000 (\frac{1}{2})^{12}\]

OpenStudy (anonymous):

But why not n-1?

OpenStudy (anonymous):

Because the formula I am given is with n-1, hmm idkk

OpenStudy (anonymous):

Do you mean the geometric formula that you are given or this question provides you a formula?

OpenStudy (anonymous):

Well that's the only formula I am given in my lesson.. yeah the first part of ur question is right

OpenStudy (anonymous):

It shouldn't be n-1. You need to visualize the situation

OpenStudy (anonymous):

Ok, then why don't you use -1? Because the first height wouldn't count right?

OpenStudy (anonymous):